OpenStudy (anonymous):

Let f(x) = -5x - 4 and g(x) = 6x - 7. Find f(x) + g(x)

OpenStudy (lochana):

f(x) + g(x) = (-5x -4) + (6x - 7) = x - 11

OpenStudy (anonymous):

thanks

OpenStudy (lochana):

bye

OpenStudy (anonymous):

bye